# MoEUT: Mixture-of-Experts Universal Transformers

### v1

@@ 1 Introduction
One important candidate is the Universal Transformer (UT, [10]). The core characteristic of UTs is recurrence in depth via sharing parameters across layers. This reintroduces the expressive power of recurrence provided by recurrent neural networks (RNNs, [11, 12, 13]). Layer sharing allows UTs to outperform regular Transformers on compositional problems such as logical inference tasks, while also yielding improvements on small-scale language modeling and translation tasks. In particular, UTs have been shown to have better compositional generalization properties [14, 15] by being able to decompose structured problems without supervision and generalize to longer sequences [16]. [^2] Dehghani et al. [10] also augment UTs with an additional adaptive computation time (ACT, [17, 18]) mechanism. However, the benefits of UTs we discuss here are purely due to layer-sharing, which, in consequence, is the focus of this work. Our models could also optionally be augmented with ACT but this is out of scope here. These empirical findings confirm that UTs are more general architectures with superior generalization properties compared to the standard Transformers, in principle. However, UTs suffer from a fundamental problem of parameter-compute ratio: sharing the parameters among $L$ layers of an $L$-layer Transformer—while keeping the same model dimensionalities—results in a model with $L$ times less parameters (ignoring the input/output layers to simply the discussion). Upscaling the size of the layer to compensate for the loss of parameters (essentially by making it $L$ times wider) usually yields a very big layer whose computational requirements in terms of compute and memory are prohibitive in practice [19, 20]. In sum, despite their potential, UTs are much less compute-efficient than standard Transformers, and thus, they are not popular for parameter-dominated tasks such as modern language modeling. Indeed, we are not aware of any previous work that has succeeded in developing compute-efficient UT models that yield competitive performance compared to standard Transformers on such tasks.
@@
Here we bring new perspectives and a solution to UTs’ fundamental compute-parameter ratio problem. We present Mixture-of-Experts Universal Transformers (MoEUTs, pronounced “moot”), a mixture-of-experts (MoE) architecture [21, 22, 23] for UTs enabling them to scale in a computationally and memory efficient way. We leverage various recent advances in MoEs for both feedforward and self-attention layers (Sec. 2.1 and 2.2), and combine them with two new innovations: (1) layer grouping, in which we recurrently stack groups of MoE-based layers, and (2) a peri-layernorm scheme (which is “in-between” the standard pre- and post-layernorm), in which we apply layer norm only before linear layers that immediately precede sigmoid or softmax activations. Both are specifically designed for shared-layer MoE architectures, and strongly supported by empirical evidence.
@@
MoEUTs allow us to build parameter- and resource-efficient UT language models outperforming standard Transformers with less compute and memory requirements on all scales on which we can afford to test (up to 1B parameters). We demonstrate its capabilities on the C4, SlimPajama, and peS2o language modeling datasets, as well as on The Stack code generation. Our experiments show that recurrence is essential for our models to achieve competitive performance. We also demonstrate good zero-shot performance on downstream tasks like BLiMP and Children’s Book Test, Lambada, HellaSwag, PIQA and ARC-E. We will provide a link to our public code repository upon acceptance.
@@ 2.2 MoE Self-Attention Layers
where ${\bm{W}}_{V}^{h,e}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$ and ${\bm{W}}_{O}^{h,e}\in\mathbb{R}^{d_{\text{head}}\times d_{\text{model}}}$ are head $h$, expert $e$, weight matrices for value and output respectively; ${\bm{s}}_{V,t}^{h}[e],{\bm{s}}_{O,t}^{h}[e]\in\mathbb{R}$ are scores of expert $e$ for head $h$ at position $t$ for value and output MoE respectively; and $\mathrm{Attention}$ denotes the standard softmax scaled dot attention [1] with ${\bm{K}}_{t}^{h}=({\bm{k}}^{h}_{1},\ldots,{\bm{k}}^{h}_{t}),{\bm{V}}_{t}^{h}=({\bm{v}}^{h}_{1},\ldots,{\bm{v}}^{h}_{t})\in\mathbb{R}^{T\times d_{\text{head}}}$ e.g., for the auto-regressive setting. Note that, here we describe position-wise computations for clarity but in practice, they can be parallelized over the tokens through matrix operations. Unlike in the original SwitcHead, which uses no regularization, we apply the same entropy regularization we use in the feedforward layer (Eq. 4) with a regularization coefficient $\delta$. (The same value is used for both value and output.)
@@ 2.3 Layer Grouping: MoE-efficient Layer Sharing & Sub-operations within an Operation
Our solution to these problems is to stack multiple layers with non-shared weights to form what we call a group of layers, reducing the number of experts in each $\sigma$-MoE while increasing the total number of attention heads. The final network is obtained by recurrently stacking such groups that share the same parameters (in a sense, redefining the group as a shared “layer” in the UT). Figure 3 provides an illustration; here, all layers denoted by “Layer A” (or “B” respectively) share the same parameters across the entire network. The size of the group, $G$, is the number of non-shared layers in it. In our experiments, the group size is between 2 and 4, and the typical number of recurrent steps is 8 or 9.
@@ 3 Main Experimental Results
Figure 4: Scaling of different models on C4 (with perplexity measured on a held-out subset of C4)
@@
We present our main experimental results on the performance and efficiency of our MoEUT on language modeling using the popular C4 dataset [39]. To demonstrate the versatility of our model, we also show our main results on the SlimPajama [40] and peS2o [41] language modeling datasets, and code generation on “The Stack” [42]. For experimental evidence in support of the benefits of shared layers for compositional generalization, we refer to much previous work (e.g., [10, 15, 14, 16, 38]). Following prior work [27, 31], we measure the compute requirements in terms of the number of multiply-accumulate (MAC) operations needed in the forward pass.
@@
Because our models are fully MoE, they decouple the number of parameters, compute and memory requirements, and different model dimensions such as $d_{\text{model}}$ and $d_{\text{ff}}$, number of layers. Thus, they provide greater flexibility for model designers. We follow a simple procedure for setting the model’s hyperparameters, as described below. All our models use RoPE positional encodings [43] with PyTorch’s fast attention implementation. The baseline models are pre-layernorm Transformers. For each baseline, we construct a parameter-matched MoEUT model. We set $d_{\text{model}}$ and the number of layers $n_{\text{layers}}$ to be the same as for the dense baseline. We use the same tokenization for each model trained on the same dataset. The number of heads $H$ for MoEUT is set to $\frac{1}{4}H$ of the corresponding dense model, $d_{\text{head}}$ is set to $2d_{\text{head}}$ of the corresponding dense model, and we set $K_{A}=2$. This matches the number of MACs spent for the value and output projections in self-attention, and reduces the number of MACs spent on calculating keys and queries and the attention matrices itself. For the $\sigma$-MoE layers, we set the expert size $d_{\text{expert}}=128$, and $K=2d_{\text{model}}/d_{\text{expert}}$. This halves the MAC requirements compared to the dense counterpart. We set the number of experts in the feedforward block, $N_{E}$, and the number of attention experts, $N_{A}$ such that the number of parameters is the same as for the dense baseline, and $10{-}15\%$ of the model’s parameter budget (excluding the embedding and classification layers) is spent in the attention computations. We set the group size $G$ to 2 for all our models below 300M parameters, $G=3$ for our 319M parameter model, and $G=4$ for the bigger models. This helps keep the number of experts manageable and improves both the performance and the speed of the model. All models are trained with batch size 64, context length 1024, for $10^{5}$ steps. This protocol allows us to perform fair comparisons between different models within our computational budget, and it leads to high quality models, as measured by our benchmarks. For more details, see Appendix A.3.
@@
To confirm the effectiveness of our model on a different task domain, here we train it on a subset of the “The Stack” dataset [42] which is a code generation task. As we cannot afford a full epoch of training, we limit ourselves to a few languages only. We use a mixture of diverse languages: Python, HTML, C++, Rust, JavaScript, Haskell, Scala, and assembly. We evaluate our models on a held-out subset of the dataset. The results are shown in Figure 6, and they are in line with our findings on the natural language domain: MoEUT outperforms the baseline.
@@
Here we evaluate the zero-shot performance of our models on six different downstream tasks: LAMBADA [44], BLiMP [45], Children’s Book Test (CBT) [46], HellaSwag [47], PIQA [48] and ARC-E [49]. For LAMBADA, we use the detokenized version from OpenAI, and we evaluate the top-1 accuracy of the last word (it can span multiple tokens; here we use greedy decoding). For CBT and BLiMP, we measure the accuracy for each task and report the average of the tasks’ accuracies. The results are shown in Table 1. We observe that our models and the baselines typically perform very similarly. MoEUT often outperforms the baseline, but the differences are marginal in all cases. This confirms that our models are indeed very capable compared to standard language models. We confirm this on peS2o and SlimPajama datasets as well.
Figure 5: Performance of MoEUT compared to a standard Transformer on The Stack.
Figure 6: Perplexity of 244M MoEUT model with different layer grouping options.
Table 1: Zero-shot downstream performance and perplexity on various language modeling datasets.
@@
Here we compare our MoEUT to another baseline, Sparse Universal Transformer (SUT; [38]) which is a recently proposed UT model that also makes use of MoE layers. We note that SUTs have not been evaluated previously on any language modeling tasks. While both MoEUT and SUT make use of an MoE for both feedforward and attention layers, there are several technical differences at various levels between the two methods: SUT uses competitive expert selection (softmax), multiple load balancing losses, and much bigger expert sizes. Their model is post-layernorm and does not use layer grouping. Unlike ours, the ACT is used in the layer dimension.
We took the original code released by Tan et al. [38] and ported it to our training pipeline for a fair comparison. As for MoEUT, we roughly match the model’s dimensionalities and number of active channels to our dense baselines. We ran a hyperparameter optimization for the regularization losses, and we found that a minimal regularization is necessary for stabilizing the training. However, larger regularization tends to hurt performance significantly. For any other hyperparameters, we set them based on their biggest translation experiments. The results are shown in Fig. 8. Effectively, SUTs, which lack our specific methods, have a significant performance disadvantage compared to our MoEUT and the parameter matched dense baseline. To the best of our knowledge, we are not aware of any prior UT architectures that are both competitive and efficient in language modeling.
@@
We investigate the effect of the layer grouping (Sec. 2.3) on our 244M parameter MoEUT model in Fig. 6. Here, $G$ denotes the number of non-shared layers within the group. $G=2$ corresponds to the model used in all other analyses. $G=1$ is a fully shared-layer model, without any grouping, and $G=18$ corresponds to the baseline fully non-shared $\sigma$-MoE model [28]. All hyperparameters are identical among all models, except for the number of MLP experts ($N_{E}$) and attention experts ($N_{A}$), which are adjusted to match the parameter count of the dense baseline. In Fig. 6, we observe that $G=2$ is optimal, and the recurrence in the layer dimension is indeed beneficial. Another interesting question is whether the grouping described in Sec. 2.3 and Fig. 3 is the right way to stack layers. Let us call the two layers in the group, A and B. The grouping we discussed so far stacks layers in the form of “ABABAB”, e.g., for a 6-layer network. An alternative is to first repeat one of the layers multiple times, followed by the repeated version of the other: “AAABBB”. The “AABB” column of Fig. 6 shows this setup for our best $G=2$ model. It can be seen that the grouping proposed in Sec. 2.3 indeed works significantly better. In fact, the AABB-style stacking is almost as bad as not doing grouping at all.
@@
Figure 7: Comparing MoEUT and SUT [38]
Figure 8: Comparing layernorm variants.
@@ 4 Analysis
Expert (re)use across layers. We first focus on whether nontrivial reuse occurs between different layers. Note that MoE-based shared-layer models could, in theory, assign different experts to different layers to “emulate” regular non-shared models. If this was the case, the model would resemble a regular Transformer instead of a Universal one. To confirm that our model is more versatile than that, we analyze whether certain experts are activated only in specific layers. We measure how many times each expert is activated in each layer. This allows us to visualize the distribution of layers that each expert prefers. To better visualize the structure, we reorder experts by a heuristic which we call “layer position score” defined as the average of the layer indices weighted by the number of expert activations in that layer. Fig. 11 shows the results. The yellow spot in the bottom right corner indicates that some experts are assigned mostly to the final layer. However, for the other experts, there is a wide range of layers where the expert is activated. Experts seem to be active in a continuous sequence of layers. This can be seen in the wide, vertically lined structure. We can conclude that MoEUT is capable of specializing in a specific layer if necessary and sharing weights between them when advantageous.
@@
Per-token expert selection diversity. Here we analyze the diversity of expert selection for a given input token across different layers and contexts. For this we measure the total number of unique experts activated for individual tokens at different layers across different positions/contexts. The result is shown in Fig. 11. On the x-axis, the tokens are ordered differently for each layer based on the number of experts used in that layer. We display the most “specialized” 1000 tokens. The minimum possible number of active experts is 16, corresponding to $K$. If only $K$ experts are consistently used by a token across different contexts, it means that the token is fully “specialized” to consistently use a single set of $K$ experts. This is almost the case for many tokens if we look at the first layer (blue curve): the number of unique experts used is low, i.e., the selection mainly depends on the token’s identity. However, in the subsequent layers, the situation is quite different: the total number of experts used increases significantly, indicating that the context of the token is taken into account for the expert selection. The diversity of the experts used peaks in the middle layers (here Layer 9) and falls slightly for layers closer to the output. For the converse analysis of expert specialization to tokens/layers, we refer to Appendix A.5.
@@
Figure 9: Instance-level average expert selection similarity between layers.
@@
Expert selection dynamics in individual columns/positions. So far, all the results have been cumulative statistics in different input sequences and positions. We might wonder about the selection behavior for the individual Transformer columns. Is expert selection mostly constant throughout the layers for individual columns of MoEUT? To answer this question, we calculate the pairwise intersection-over-union of the set of selected experts between all layers in individual columns and average this metric over the whole validation set. We show the result in Fig. 9. There is a non-negligible overlap between the selected experts in subsequent layers; however, it is far from complete overlap. This indicates that experts usually change dynamically in a single column, performing different functionality in different layers.
@@
Figure 10: Layer preference of different experts.
Figure 11: No. of experts used in different layers.

### v2

@@ 1 Introduction
One important candidate is the Universal Transformer (UT, [10]). The core characteristic of UTs is recurrence in depth via sharing parameters across layers. This reintroduces the expressive power of recurrence provided by recurrent neural networks (RNNs, [11, 12, 13]). Layer sharing allows UTs to outperform regular Transformers on compositional problems such as logical inference tasks, while also yielding improvements on small-scale language modeling and translation tasks. In particular, UTs have been shown to have better compositional generalization properties [14, 15] by being able to decompose structured problems without supervision and generalize to longer sequences [16]. [^2] Dehghani et al. [10] also augment UTs with an additional adaptive computation time (ACT, [17, 18]) mechanism. However, the benefits of UTs we discuss here are purely due to layer-sharing, which, in consequence, is the focus of this work. Our models could also optionally be augmented with ACT but this is out of scope here. These empirical findings confirm that UTs are more general architectures with superior generalization properties compared to standard Transformers, in principle. However, UTs suffer from a fundamental problem of parameter–compute ratio: sharing the parameters among $L$ layers of an $L$-layer Transformer—while keeping the same model dimensionalities—results in a model with $L$ times fewer parameters (ignoring the input/output layers to simplify the discussion). Upscaling the size of the layer to compensate for the loss of parameters (essentially by making it $L$ times wider) usually yields a very big layer whose computational requirements in terms of compute and memory are prohibitive in practice [19, 20]. In sum, despite their potential, UTs are much less compute-efficient than standard Transformers, and thus, they are not popular for parameter-dominated tasks such as modern language modeling. Indeed, we are not aware of any previous work that has succeeded in developing compute-efficient UT models that yield competitive performance compared to standard Transformers on such tasks.
@@
Here we bring new perspectives and a solution to UTs’ fundamental compute–parameter ratio problem. We present Mixture-of-Experts Universal Transformers (MoEUTs, pronounced “moot”), a mixture-of-experts (MoE) architecture [21, 22, 23] for UTs enabling them to scale in a computationally and memory efficient way. We leverage various recent advances in MoEs for both feedforward and self-attention layers (Sec. 2.1 and 2.2), and combine them with two new innovations: (1) layer grouping, in which we recurrently stack groups of MoE-based layers, and (2) a peri-layernorm scheme (which is “in-between” the standard pre- and post-layernorm), in which we apply layer norm only before linear layers that immediately precede sigmoid or softmax activations. Both are specifically designed for shared-layer MoE architectures, and strongly supported by empirical evidence.
@@
MoEUTs allow us to build parameter- and resource-efficient UT language models outperforming standard Transformers with less compute and memory requirements on all scales on which we can afford to test (up to 1B parameters). We demonstrate their capabilities on the C4, SlimPajama, and peS2o language modeling datasets, as well as on The Stack code generation. Our experiments show that recurrence is essential for our models to achieve competitive performance. We also demonstrate good zero-shot performance on downstream tasks like BLiMP and Children’s Book Test, Lambada, HellaSwag, PIQA and ARC-E.
@@ 2.2 MoE Self-Attention Layers
where ${\bm{W}}_{V}^{h,e}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$ and ${\bm{W}}_{O}^{h,e}\in\mathbb{R}^{d_{\text{head}}\times d_{\text{model}}}$ are head $h$, expert $e$, weight matrices for value and output respectively; ${\bm{s}}_{V,t}^{h}[e],{\bm{s}}_{O,t}^{h}[e]\in\mathbb{R}$ are scores of expert $e$ for head $h$ at position $t$ for value and output MoE respectively; and $\mathrm{Attention}$ denotes the standard softmax scaled dot attention [1] with ${\bm{K}}_{t}^{h}=({\bm{k}}^{h}_{1},\ldots,{\bm{k}}^{h}_{t}),{\bm{V}}_{t}^{h}=({\bm{v}}^{h}_{1},\ldots,{\bm{v}}^{h}_{t})\in\mathbb{R}^{T\times d_{\text{head}}}$ e.g., for the auto-regressive setting. Note that, here we describe position-wise computations for clarity but in practice, they can be parallelized over the tokens through matrix operations. Unlike in the original SwitchHead, which uses no regularization, we apply the same entropy regularization we use in the feedforward layer (Eq. 4) with a regularization coefficient $\delta$. (The same value is used for both value and output.)
@@ 2.3 Layer Grouping: MoE-efficient Layer Sharing & Sub-operations within an Operation
Our solution to these problems is to stack multiple layers with non-shared weights to form what we call a group of layers, reducing the number of experts in each $\sigma$-MoE while increasing the total number of attention heads. The final network is obtained by recurrently stacking such groups that share the same parameters (in a sense, redefining the group as a shared “layer” in the UT). Fig. 3 provides an illustration; here, all layers denoted by “Layer A” (or “B” respectively) share the same parameters across the entire network. The size of the group, $G$, is the number of non-shared layers in it. In our experiments, the group size is between 2 and 4, and the typical number of recurrent steps is 8 or 9.
@@ 3 Main Experimental Results
Figure 4: Scaling of different models on C4 (with perplexity measured on a held-out subset of C4). (a) MoEUT slightly outperforms parameter-matched models with no layer sharing. The gap grows with scale. (b) Given equal amounts of compute, MoEUT outperforms other models by a large margin.
@@
We present our main experimental results on the performance and efficiency of MoEUT on language modeling using the popular C4 dataset [39]. To demonstrate the versatility of our model, we also show our main results on the SlimPajama [40] and peS2o [41] language modeling datasets, and code generation on “The Stack” [42]. For experimental evidence in support of the benefits of shared layers for compositional generalization, we refer to much previous work (e.g., [10, 15, 14, 16, 38]). Following prior work [27, 31], we measure the compute requirements in terms of the number of multiply-accumulate (MAC) operations needed in the forward pass.
@@
Because our models are fully MoE, they decouple the number of parameters, compute and memory requirements, and different model dimensions such as $d_{\text{model}}$ and $d_{\text{ff}}$, number of layers. Thus, they provide greater flexibility for model designers. We follow a simple procedure for setting the model’s hyperparameters, as described below. All our models use RoPE positional encodings [43] with PyTorch’s fast attention implementation. The baseline models are pre-layernorm Transformers. For each baseline, we construct a parameter-matched MoEUT model. We set $d_{\text{model}}$ and the number of layers $n_{\text{layers}}$ to be the same as for the dense baseline. We use the same tokenization for each model trained on the same dataset. The number of heads $H$ for MoEUT is set to $\frac{1}{4}H$ of the corresponding dense model, $d_{\text{head}}$ is set to $2d_{\text{head}}$ of the corresponding dense model, and we set $K_{A}=2$. This matches the number of MACs spent for the value and output projections in self-attention, and reduces the number of MACs spent on calculating keys and queries and the attention matrices itself. For the $\sigma$-MoE layers, we set the expert size $d_{\text{expert}}=128$, and $K=2d_{\text{model}}/d_{\text{expert}}$. This halves the MAC requirements compared to the dense counterpart. We set the number of experts in the feedforward block, $N_{E}$, and the number of attention experts, $N_{A}$ such that the number of parameters is the same as for the dense baseline, and $10{-}15\%$ of the model’s parameter budget (excluding the embedding and classification layers) is spent in the attention computations. We set the group size $G$ to 2 for all our models below 300M parameters, $G=3$ for our 319M parameter model, and $G=4$ for the bigger models. This helps keep the number of experts manageable and improves both the performance and the speed of the model. All models are trained with batch size 64 and context length 1024, for $10^{5}$ steps. This protocol allows us to perform fair comparisons between different models within our computational budget, and it leads to high quality models, as measured by our benchmarks. For more details, see Appendix A.4.
@@
To confirm the effectiveness of our model on a different task domain, here we train it on a subset of the “The Stack” dataset [42] which is a code generation task. As we cannot afford a full epoch of training, we limit ourselves to a few languages only. We use a mixture of diverse languages: Python, HTML, C++, Rust, JavaScript, Haskell, Scala, and assembly. We evaluate our models on a held-out subset of the dataset. The results are shown in Fig. 6, and they are in line with our findings on the natural language domain: MoEUT outperforms the baseline.
@@
Here we evaluate the zero-shot performance of our models on six different downstream tasks: LAMBADA [44], BLiMP [45], Children’s Book Test (CBT) [46], HellaSwag [47], PIQA [48], and ARC-E [49]. For LAMBADA, we use the detokenized version from OpenAI, and we evaluate the top-1 accuracy of the last word (it can span multiple tokens; here we use greedy decoding). For CBT and BLiMP, we measure the accuracy for each task and report the average of the tasks’ accuracies. The results are shown in Tab. 1. We observe that our models and the baselines typically perform very similarly. MoEUT often outperforms the baseline, but the differences are marginal in all cases. This confirms that our models are indeed very capable compared to standard language models. We confirm this on peS2o and SlimPajama as well.
Figure 5: Performance of MoEUT compared to a standard Transformer on The Stack. MoEUT outperforms standard Transformers. The gap grows with scale.
Figure 6: Perplexity of 244M MoEUT models with different layer grouping options. A small group size of $G=2$ works the best, showing the advantage of layer sharing.
Table 1: Zero-shot downstream performance and perplexity on various language modeling datasets. MoEUT marginally outperforms standard Transformers in most tasks, confirming that MoEUT is indeed a capable language model.
@@
Here we compare our MoEUT to another baseline, Sparse Universal Transformer (SUT; [38]), which is a recently proposed UT model that also makes use of MoE layers. We note that SUTs have not been evaluated previously on standard language modeling tasks. While both MoEUT and SUT make use of an MoE for both feedforward and attention layers, there are several technical differences at various levels between the two methods: SUT uses competitive expert selection (softmax), multiple load balancing losses, and much bigger expert sizes. Their model is post-layernorm and does not use layer grouping. Unlike ours, Adaptive Computation Time (ACT) is used in the layer dimension.
We took the original code released by Tan et al. [38] and ported it to our training pipeline for a fair comparison. As for MoEUT, we roughly match the model’s dimensionalities and number of active channels to our dense baselines. We ran a hyperparameter optimization for the regularization losses, and we found that a minimal regularization is necessary for stabilizing the training. However, larger regularization tends to hurt performance significantly. All other hyperparameters are set based on Tan et al.’s biggest translation experiments. The results are shown in Fig. 8. Effectively, SUTs, which lack our specific methods, have a significant performance disadvantage compared to our MoEUT and the parameter-matched dense baseline. Upon careful investigation, we found that most of this poor performance comes from the ACT mechanism that the authors advertise as one of the main components of their model. After removing the ACT, the performance improves dramatically. However, even with this setup, it underperforms both MoEUT and the standard Transformer baseline. This is also confirmed on downstream tasks in Tab. 2 in the appendix. Moreover, as we show in Appendix A.7, our model runs much faster and uses only a fraction of the memory required for the SUT. To the best of our knowledge, we are not aware of any prior UT architectures that are both competitive and efficient in language modeling.
@@
We investigate the effect of the layer grouping (Sec. 2.3) on our 244M parameter MoEUT model in Fig. 6. Here, $G$ denotes the number of non-shared layers within the group. $G=2$ corresponds to the model used in all other analyses. $G=1$ is a fully shared-layer model, without any grouping, and $G=18$ corresponds to the baseline fully non-shared $\sigma$-MoE model [28]. All hyperparameters are identical among all models, except for the number of MLP experts ($N_{E}$) and attention experts ($N_{A}$), which are adjusted to match the parameter count of the dense baseline. In Fig. 6, we observe that $G=2$ is optimal, and the recurrence in the layer dimension is indeed beneficial. Another interesting question is whether the grouping described in Sec. 2.3 and Fig. 3 is the right way to stack layers. Let us call the two layers in the group A and B. The grouping we discussed so far stacks layers in the form of “ABABAB”, e.g., for a 6-layer network. An alternative is to first repeat one of the layers multiple times, followed by the repeated version of the other: “AAABBB”. The “AABB” column of Fig. 6 shows this setup for our best $G=2$ model. It can be seen that the grouping proposed in Sec. 2.3 indeed works significantly better. In fact, the AABB-style stacking is almost as bad as not doing grouping at all.
@@
Figure 7: Comparing MoEUT and SUT [38]. MoEUT outperforms both the original SUT and our improved version by a large margin.
Figure 8: Comparing layernorm variants. Peri-layernorm outperforms both pre- and post-layernorm.
@@ 4 Analysis
Expert (re)use across layers. We first focus on whether nontrivial reuse occurs between different layers. Note that MoE-based shared-layer models could, in theory, assign different experts to different layers to “emulate” regular non-shared models. If this were the case, the model would resemble a regular Transformer instead of a Universal one. To confirm that our model is more versatile than that, we analyze whether certain experts in the MLP layers are activated only in specific layers. We measure how many times each expert is activated in each layer. This allows us to visualize the distribution of layers that each expert prefers. To better visualize the structure, we reorder experts by a heuristic which we call “layer position score” defined as the average of the layer indices weighted by the number of expert activations in that layer. Fig. 10 shows the results. The yellow spot in the bottom right corner indicates that some experts are assigned mostly to the final layer. However, for the other experts, there is a wide range of layers where the expert is activated. Experts seem to be active in a continuous sequence of layers. This can be seen in the wide, vertically lined structure. We can conclude that MoEUT is capable of specializing in a specific layer if necessary and sharing weights between them when advantageous.
@@
Per-token expert selection diversity. Here we analyze the diversity of expert selection in the MLP layers for a given input token across different layers and contexts. For this we measure the total number of unique experts activated for individual tokens at different layers across different positions/contexts. The result is shown in Fig. 10. On the x-axis, the tokens are ordered differently for each layer based on the number of experts used in that layer. We display the most “specialized” 1000 tokens. The minimum possible number of active experts is 16, corresponding to $K$. If only $K$ experts are consistently used by a token across different contexts, it means that the token is fully “specialized” to consistently use a single set of $K$ experts. This is almost the case for many tokens if we look at the first layer (blue curve): the number of unique experts used is low, i.e., the selection mainly depends on the token’s identity. However, in the subsequent layers, the situation is quite different: the total number of experts used increases significantly, indicating that the context of the token is taken into account for the expert selection. The diversity of the experts used peaks in the middle layers (here Layer 9) and falls slightly for layers closer to the output. For the converse analysis of expert specialization to tokens/layers, we refer to Appendix A.6.
Expert selection dynamics in individual columns/positions. So far, all the results have been cumulative statistics in different input sequences and positions. We might wonder about the selection behavior for the individual Transformer columns. [^3] By representing the Transformer’s activations in a 2D grid, with token positions on the x-axis and depth on the y-axis, “columns” correspond to all hidden activations across depth given a token position. Is expert selection mostly constant throughout the layers for individual columns of MoEUT?
Figure 9: Layer preference of different experts. Most experts are used in multiple layers, while some of them (see, bottom right) specialize to certain layers, showing the flexibility of our model.
Figure 10: No. of unique experts used in different layers. Tokens are routed to many different experts (depending on the context), especially in the middle layers.
@@
Figure 11: Instance-level average expert selection similarity between layers. Individual tokens are routed to a diverse set of experts across the layers.
@@
To answer this question, we calculate the pairwise intersection-over-union of the set of selected experts between all layers in individual columns and average this metric over the whole validation set. We show the result in Fig. 11. There is a non-negligible overlap between the selected experts in subsequent layers; however, it is far from complete overlap. This indicates that experts usually change dynamically in a single column, performing different functionality in different layers.